a coin is dropped from a height of 750 feet. the height,s,at time,t, is given by s=-16t^2+750.
a. find the average velocity on the interval[1,3]
b. find the instantaneous velocity when t=3
c. find the velocity of the coin when it hits the ground

Answer by stanbon(75887)

a coin is dropped from a height of 750 feet. the height,s,at time,t, is given by s=-16t^2+750.
a. find the average velocity on the interval[1,3]
Average = [s(3)-s(1)]/[3-1}
----------
b. find the instantaneous velocity when t=3
Derivative = -32t
Instantaneous velocity at t=3 = -32(3) = -96 ft/sec
---------------
c. find the velocity of the coin when it hits the ground
When it hits the ground s(t) = 0
Solve -16t^2+750 = 0
t^2 = 750/16
t = 6.847 seconds
Velocity at t=6.847 sec = -32(6.847)= -219.09 ft/sec
=========
Cheers,
Stan H.